The reaction involved in the explosive combusion of acetylene is : 2 C2H2 (g) + 5 O2(g) g 4 CO2(g) + 2 H2 (g) How many L of CO2 gas (measured at STP) will be formed during the combustion of 50.0 L C2H2 gas (measured at STP)


1
Expert's answer
2021-11-05T01:07:41-0400

2C2H2(g) + 5O2(g) g → 4CO2(g) + 2H2(g)

At STP one mole of any gas occupies 22.4 L

n(C2H2) "= \frac{50.0}{22.4} = 2.232 \;mol"

According to the reaction:

n(CO2) = 2n(C2H2) "= 2 \times 2.232 = 4.464 \;mol"

V(CO2) "= 4.464 \times 22.4 = 99.99 ≈ 100 \;L"

Answer: 100 L